find the volume of the triangular prism..
the base is a right angled triangle. The hypotenuse is 44cm and angle is 52 between the hypotenuse and the adjacent, the height of the prism is 56 cm..
the answer at the back is 26 299cm3 to the nearest cm... but i can't seem to get this answer any how... solution please...

Answer by edjones(8007)

a/sin A = b/sin B (Law of sines)
44/sin 90 = b/sin 52
44=b/.788
b=34.67 cm
.
a/sin A = c/sin C
44/sin 90 = c/sin 38
44=c/.616
c=27.09 cm
.
V=.5abh
=.5*34.67*27.09*56
=26,299 cm^3
.
Ed